Title

06:03 ESM Preset
  • RunAO Failed
  • Skip Frame
  • Actuator 557 reporting 158 um

06:05 PresetAO from Command GUI

06:10 Camera Lens not moving
  • Power Camera lens
  • Stop/Start camera lens process

06:15 Send new Script

06:19 Set Intervention Mode

06:23 Power Off WUnit
  • stop/start processes
  • Power On

07:00 Alfio online
  • Tried ACE Preset, still get skipfrimes
  • Alfio finds the the Flat has many actuators with positions around 46 um. Limit is 44 um. May be getting skip frame from position.

07:08 ACE Preset
  • PresetAO Success
  • RunAO Success
    • bin 2
    • 400 Hz
    • 153
  • Many Offsets success
  • Offset outside of FOV. remain paused and resume when back in FOV. Success
  • Two problems (we found)
    1. Poor flat applied: Few modes ~40 and Tip on flat. This limited our ability to correct large tip/tilt. Get many skip frames
      • he current flat has 48 modes...
    2. For ESM our table still has new, bad reconstructors (KL_v21) which has problems with Tip/Tilt. Get many skip frames.
  • Now that we are in full ACE AO, using old (better) KL_v16 reconstructors and not having T/T problems
  • With AO loop closed with 153 modes, then flat of 48 is now corrected up to 153

07:54 Optical Loop Data
  • 075419
  • 080116
  • 080506
  • 080659
  • Alfio, Enrico and Simone look at data and looks good
    • Think maybe should change mode for splitting of HO1 and HO2 gains
    • slit should be at ~100 for both Bin 1 and Bin2
  • In Luci images see astigmatism (cross) Need to apply NCPA for bin 2 for LUCI2 (poor metal optics)

08:37 ACE Preset (Will turn on NCPA) (same object at before)
  • Put Pause after preset in LUCI script so we can turn on NCAP
  • No GOPT data set
    • 084307
  • Gopt On but no NCPA
    • Opt Gain = 0.15
    • 084459
    • 084518
  • NCPA On
    • 084611
    • 084625

08:49 Collect Science Data with offsets
  • 084930
  • 085108
  • Optical gain down to 0.03
    • Maybe because of many quick offset
    • Finally get a long image with many Gopt calculation so go back to 0.13
  • Opt Gain = 0.13
    • 090022
  • Opt Gain = 0.2
    • 090554
    • 091319
    • 091948
    • 092338
  • Opt Gain = 0.16
    • 092934

09:58 RunAO Failed
  • Measure Ref Star mag < 5 mag from sent mag
  • Sent 9.5
  • Measure 14.7

10:03 ACE Preset with AO Ref man 14.
  • Measure 14.2
  • Start bin 3, switch to bin 4
  • did not turn on NCPA
  • Bin 4
  • 312 Hz
  • 36 modes

10:12 Opt Loop Data, no NCPA
  • 101224
  • 101359
  • 101526

10:42 ACE Preset
  • Bin 3
  • 200 Hz
  • 66 modes
  • Opt loop data
    • 104238
    • 104404
    • 104750
    • 104950
    • 105115
    • 105639
    • 110423 Offet about half way though saving set

11:29 ACE Preset (to PSF star)
  • measured 13.8 mag, same AO parameters
  • Bin 3
  • 66 modes
  • 312 Hz

11:30 Stop observing as background becomes too bright

-- %USERSIG{DougMiller - 2017-06-14}%
Topic revision: r2 - 14 Jun 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback